# Boosting Cardiac Color Doppler Frame Rates with Deep Learning

### v1

@@ 1 Introduction
Color Doppler echocardiography facilitates the concurrent visualization of cardiac tissue and intracardiac blood movement. In clinical settings, this imaging modality may serve as a visual aid to diagnose some cardiac pathologies. However, the frame rate currently available is insufficient to quantitatively assess blood velocities during a cardiac cycle. Increasing the temporal resolution of color Doppler could potentially open up new possibilities for diagnosing diastolic impairment through the use of quantitative analysis tools [1].
@@
    We generated a training dataset for supervised learning by simulating color Doppler echocardiography with a dedicated pipeline. It included customized data augmentation techniques to increase the dataset size.
@@
    We implemented state-of-the-art deep learning architectures for Doppler velocity estimation from I/Q signals with a reduced packet size. We explored the performance of U-Net and ConvNeXt architectures, and the impact of using real-valued versus complex-valued representations.
@@
    We validated the proposed deep learning models using both in vitro and in vivo sequences. These experiments showed that the deep learning models achieved comparable results to the autocorrelator while exhibiting robustness to noise and inherent reduction of aliasing.
@@ 2.2 Color Doppler simulations
Training supervised deep learning models requires large annotated datasets. In the context of color Doppler, high-quality in vivo ground truth is not available. Therefore, it is necessary to rely on simulations to build a supervised training dataset. In this study, we adapted the patient-based color Doppler simulator developed by Sun et al. [15]. A schematic representation of the inputs and outputs of the simulation pipeline is illustrated in Figure 1.
@@ 2.2.2 Modeling blood flow motion
Patient Doppler velocity frames were treated as independent due to the low frame rate of color Doppler. For each color Doppler frame, a 2-D intraventricular vector flow map was computed from the patient’s Doppler velocities using a physics-based method [17]. For each frame, blood scatterers were distributed in the intraventricular region. Series of $[\text{number of firings}\times n]$ time steps were defined, during which the blood scatterers were displaced according to the 2-D velocity maps. To enhance realism, fluctuation values were computed from velocity maps and added to velocity values. This process was repeated for all frames of the sequence. From the resulting vector outputs, we used the radial components to build the ground truth Doppler velocity maps $D\in\mathbb{R}^{h\times w}$ in our training dataset.
@@ 2.2.3 Simulation of an ultrasound acquisition
From the tissue and blood scatterer position maps defined at each time step, scanline-based ultrasound acquisitions were simulated using SIMUS [18, 19] of the Matlab UltraSound Toolbox (MUST) [20]. The RF signals were post-processed conventionally with I/Q-demodulation and delay-and-sum beamforming on a $h\times w$ grid. The resulting I/Q signals $S\in\mathbb{C}^{h\times w\times n}$ were the inputs in our training dataset.
@@ 2.3.1 U-Net
  • Real U-Net. The input of the network was in the form $S\in\mathbb{R}^{h\times w\times 2n}$, where the real and imaginary parts of the I/Q signals were concatenated to obtain real representations. We began with the standard U-Net architecture as described in [21] and made two modifications. Firstly, to ensure a lightweight network, we retained two dimension reduction steps, thereby reducing the network to only 6 layers. Secondly, the kernels of the two convolutions at each U-Net stage were of sizes $5\times 5$ and $3\times 3$. The number of feature maps was $(32,64,128)$ for the encoder part and symmetrically for the decoder part. With this configuration, the network had 1.5 million parameters, and the lowest dimension achieved was 45 $\times$ 10.
@@ 2.3.2 ConvNeXt
We then considered ConvNeXt [24], a recent convolutional neural network architecture. Since 2020, hierarchical vision Transformer architectures have demonstrated superior performance over convolutional models in various computer vision tasks. However, Transformers models suffer from being data-intensive. As a solution, ConvNeXt was introduced in 2022. It is a convolutional-based model that integrates strategies from Transformer models, resulting in an architecture that outperforms state-of-the-art Transformers while remaining data efficient. One of its key features is the separate handling of spatial and channel information, akin to the self-attention mechanism in Transformers. This property is achieved through the use of depthwise convolutions. We adapted the ConvNeXt architecture to our problem by utilizing a U-Net architecture with ConvNeXt as its backbone (i.e., ConvNeXt layers served as feature extractors in the encoder part). A schematic representation of the considered architecture is depicted in Figure 2.
@@ 2.4.1 Zoomed samples
[…]
@@ 2.4.2 Aliased samples
Aliasing patterns appear in Doppler velocity maps when the maximum speed exceeds the Nyquist velocity. In such cases, phase unwrapping methods can be applied to the Doppler velocity maps as a post-processing step, as described in the introduction. In contrast to these post-processing approaches, we aimed to make the Doppler estimation inherently robust to aliasing. This was achieved by providing the deep learning models with new training pairs $(S,D)$, where $S$ represents the I/Q aliased signal and $D$ denotes the corresponding alias-free Doppler ground truth. The aliased samples were obtained during simulations by randomly reducing the original PRF values of each sequence by a factor in the $[0.4,0.6]$ range. Each simulated aliased sample was considered as an augmented sample and added to the dataset with a probability of 0.5.
@@ 3.1.1 In silico experiment
PRF patient-based
sector width patient-based
depth patient-based
@@
During the simulations, we leveraged patient-specific parameters as pulse repetition frequency (PRF), sector angle, and depth. The simulated RF signals were demodulated into I/Q signals and beamformed into a $180\times 40$ grid using delay-and-sum. Each training pair $(S,D)$ consisted of I/Q signals $S\in\mathbb{C}^{180\times 40\times n}$ and Doppler velocity references $D\in\mathbb{R}^{180\times 40}$. Figure 3A shows a simulated sample.
Figure 3: Example of a simulated echocardiography acquisition with its original (A) and decreased (B) PRF. The Doppler velocities were estimated using an autocorrelator in both cases.
Data augmentation, as described in Section 2.4.1, was applied. Samples with reduced packet size were created by selecting two consecutive frames in the slow-time. The augmented dataset of color Doppler simulations contained $\sim~{}4,500$ pairs and is referred to as Alias-free-Set from here on.
Only 24 of the original samples contained aliased regions, accounting for less than 0.001% of the total dataset pixels. As a consequence, we used the procedure described in Section 2.4.2. The subset of aliased samples is referred to as Aliased-Set in the sequel. An example of a sample simulated with both its original and decreased PRF is depicted in Figure 3. The Alias-free-Set dataset augmented with the Aliased-Set contained $\sim~{}5,300$ pairs, henceforth referred to as Full-Set.
@@ 3.1.2 In vitro experiment
We designed a rotating disk experiment to measure Doppler velocities from a color Doppler acquisition. We constructed a 6 cm diameter agar disk and connected it to a rotating motor. A Verasonics P4-2v cardiac probe was positioned at 8 cm from the center of the disk. To minimize clutter and artifacts, we submerged the entire setup in a water tank. Figure 4A provides a schematic representation of the experiment.
Figure 4: (A) Schematic representation of the in vitro experiment using a rotating disk phantom. (B) B-mode obtained during one of the experiments.
The acquisition parameters were set as described in Table 1, with the sector width, depth and PRF configured to $50^{\circ}$, 12 cm and 6,000 Hz, respectively. The packet size was $n=32$. We ran a total of 12 experiments, varying the maximum outer speed between 0.1 and 1.13 m/s, mirroring intracardiac blood velocities. The Nyquist velocity was $v_{N}=0.85$ m/s for all experiments, resulting in four experiments displaying aliased regions. An example of an acquisition result is depicted in Figure 4B.
@@ 3.1.3 In vivo experiment
Cardiac echo-Doppler focused acquisitions were conducted on two healthy volunteers. The study received approval from the ethics and research committee of Polytechnique Montréal (CER-2122-54-D). In accordance with the simulation and phantom experiments, the P4-2v Verasonics probe was used with identical parameters as those used for the phantom. Data were acquired in the apical three-chamber view, including both B-mode and Doppler information. The packet size was set to $n=16$. The ultrasound signals were post-processed and beamformed on a $180\times 40$ grid. We obtained 16 color Doppler image samples from each volunteer.
[…]
Samples with reduced packet size were generated by selecting two consecutive frames in the slow-time. In the in vivo experiments, ground truth Doppler velocity maps were not available. Hence, Doppler velocity maps obtained with the autocorrelator using the maximum packet size ($n=16$) were used as references.
@@ 3.2.1 Training settings
Because the ultrasound data were simulated with patient-specific PRFs, the Nyquist velocity varied from sample to sample. Therefore, we opted to use phase shifts as the ground truth instead of Doppler velocities to ensure standardized inputs across samples. Each sample was normalized with respect to its maximum complex modulus across the slow-time. For greater robustness, we trained the network using 9-fold cross-validation, where samples from a given patient were assigned to the same fold. Once the test fold was chosen, the split of samples in training/validation was set to 9/1. The network weights were initialized with the Xavier distribution. The loss function was the masked mean squared error, where the mask corresponded to the left ventricular cavity. The optimizer was AdamW, and the batch size was set to 16. The initial value of the learning rate was set to $10^{-3}$ and decreased following a plateau scheduler with a patience of 10 epochs. Trainings were performed using the PyTorch library on an NVIDIA V100 GPU with 16 GB of memory.
@@ 4.1 In silico results
The performance of our methods in Doppler velocity estimation, when trained on Alias-free-Set and using a packet size $n=2$, is presented in Table 2. For all methods, the metrics were computed exclusively within the left ventricle.
[…]
Method Alias-free-Set
Autocorrelator $3.7\pm 2.6$
Real U-Net $1.8\pm 0.9$
Complex U-Net $1.8\pm 0.8$
ConvNeXt $\boldsymbol{1.7\pm 0.9}$
@@
We examined the samples with higher errors for both the autocorrelator and the deep learning models. For the former, large estimation errors mainly occurred in two situations: noisy signals and in the presence of aliasing. For the deep learning models, large errors were observed in two specific settings: in a few pixels near the valves and the presence of aliasing across a large region of the ventricle. To mitigate the effects of aliasing, we trained our models on the Full-Set dataset. The results of the estimation are presented in Table 3.
[…]
Method Alias-free-Set Aliased-Set
Autocorrelator $3.7\pm 2.6$ $16\pm 9.8$
Real U-Net $1.9\pm 0.8$ $2.5\pm 1.4$
Complex U-Net $1.8\pm 0.8$ $2.4\pm 1.6$
ConvNeXt $\boldsymbol{1.7\pm 0.7}$ $\boldsymbol{2.3\pm 1.5}$
We noted that when testing on Alias-free-Set, we maintained a consistent performance compared to the models trained on Alias-free-Set. Moreover, when testing on Aliased-Set, we observed a substantial reduction in error compared to the autocorrelator method, indicating that the models are capable to alleviate aliasing to some extent.
@@
Figure 5: Doppler velocity estimations for five in silico samples. B-mode and color Doppler images were simulated from the data of five different patients. Samples A and B had velocities below the Nyquist limit. Samples C, D and E had velocities above the Nyquist limit. Aliased regions are indicated with white arrows.
We only report predictions obtained with ConvNeXt (the best-performing model) for the sake of brevity. As explained above, ConvNeXt trained in both ways achieved accurate estimates in the absence of aliasing (samples A and B). In such cases, we observed that the autocorrelator was more susceptible to noise. When encoutering small aliased regions (sample C, left region), the model trained with Alias-free-Set effectively removed aliasing. However, when aliasing extended over a large region, it either smoothed the region, leading to velocity underestimation (sample C, right part, and sample D), or produced a pattern similar to that of the autocorrelator, albeit smoothed at the edges (sample E). In contrast, ConvNeXt trained on Full-Set yielded results closely aligned with the ground truth.

### v2

@@ 1 Introduction
Color Doppler echocardiography facilitates the concurrent visualization of cardiac tissue and intracardiac blood movement. In clinical settings, this imaging modality is extensively used as a visual aid to assess cardiac function such as valvular regurgitation, ventricular filling and ventricular diastolic function [1]. The frame rate achieved by focused Color Doppler echocardiography is in the range of 10-30 fps assuming an apical long-axis acquisition from the apex to the basis of 12 cm and a number of lines between 20 and 60, depending on the sector width. Considering a standard subject with 80 bpm, this results in 8 to 23 frames per cardiac cycle. However, Krovetz et al. [2] found that around 25 harmonics are needed to accurately retrieve blood flow accelerations within the left ventricle, indicating that a minimum frame rate of 25 frames per cardiac cycle is necessary to quantitatively assess blood motion during all the phases of the cardiac cycle. Therefore, increasing the temporal resolution of color Doppler could allow a better characterization of intracardiac blood movement and potentially open up new diagnosis possibilities. In particular, Faurie et al. [3] investigated the use of ultrafast imaging to achieve high-frame rates of around 80 fps for color Doppler imaging. They found that high frame rates allowed to study blood vortex dynamics, and they derived quantitative tools that could be used to assess diastolic impairment. Here, we were interested in achieving high-frame rate color Doppler in the conventional framework of focused imaging.
@@
    We generated a training dataset for supervised learning by simulating color Doppler echocardiography with a dedicated pipeline. It included the generation of high-quality Doppler velocity ground truths and the simulation of realistic acquisitions of ultrasound signals.
@@
    We adapted state-of-the-art deep learning architectures for real-time Doppler velocity estimation from I/Q signals with a reduced packet size. We explored the performance of U-Net and ConvNeXt architectures, and the impact of using real-valued versus complex-valued representations.
@@
    We introduced a tailored zooming data augmentation procedure to increase the number of samples while ensuring signal coherence. In addition, we generated augmented aliased samples to make the models inherently robust to aliasing.
  4. 4.
    We validated the proposed deep learning models using both in vitro and in vivo sequences. These experiments showed that the deep learning models were able to generalize to domain shift data. They achieved comparable results to the autocorrelator while exhibiting robustness to noise and inherent reduction of aliasing.
@@ 2.2 Color Doppler simulations
Training supervised deep learning models requires large annotated datasets. In the context of color Doppler, high-quality in vivo ground truth is not available. Using directly the patient Doppler data, the models would learn to reproduce the behavior of the standard estimator along with its inaccuracies. Therefore, we decided to rely on a simulation approach that allows to generate a supervised training dataset with reliable ground truths. In this study, we adapted the patient-based color Doppler simulator developed by Sun et al. [17]. Unless otherwise specified, the parameters were consistent with those described in the original paper. A schematic representation of the inputs and outputs of the simulation pipeline is illustrated in Figure 1.
@@ 2.2.2 Modeling blood flow motion
Patient Doppler velocity frames were treated as independent due to the low frame rate of color Doppler. For each color Doppler frame, a 2-D intraventricular vector flow map was computed from the patient’s Doppler velocities using a physics-based method [19]. The method was used as described in the original paper. Then, for each frame, blood scatterers were distributed in the intraventricular region. Series of $[\text{number of firings}\times n]$ time steps were defined, during which the blood scatterers were displaced according to the 2-D velocity maps. To enhance realism, fluctuation values were computed from velocity maps and added to velocity values. This process was repeated for all frames of the sequence. From the resulting vector outputs, we used the radial components to build the ground truth Doppler velocity maps $D\in\mathbb{R}^{h\times w}$ in our training dataset.
@@ 2.2.3 Simulation of an ultrasound acquisition
From the tissue and blood scatterer position maps defined at each time step, scanline-based ultrasound acquisitions were simulated using SIMUS [20, 21] of the Matlab UltraSound Toolbox (MUST) [22]. The RF signals were post-processed conventionally with I/Q-demodulation and delay-and-sum beamforming on a $h\times w$ grid. The resulting clutter-free beamformed I/Q signals $S\in\mathbb{C}^{h\times w\times n}$ were the inputs in our training dataset.
@@ 2.3.1 U-Net
  • Real U-Net. The input of the network was in the form $S\in\mathbb{R}^{h\times w\times 2n}$, where the real and imaginary parts of the I/Q signals were concatenated to obtain real representations. We began with the standard U-Net architecture as described in [23] and made two modifications. Firstly, to ensure a lightweight network, we retained two dimension reduction steps, thereby reducing the network to only 6 layers. Secondly, the kernels of the two convolutions at each U-Net stage were of sizes $5\times 5$ and $3\times 3$. The number of feature maps was $(32,64,128)$ for the encoder part and symmetrically for the decoder part. A scheme of the architecture is shown in Figure 2. With this configuration, the network had 1.5 million parameters, and the lowest dimension achieved was 45 $\times$ 10.
@@ 2.3.2 ConvNeXt
We then considered ConvNeXt [26], a recent convolutional neural network architecture. Since 2020, hierarchical vision Transformer architectures have demonstrated superior performance over convolutional models in various computer vision tasks. However, Transformers models suffer from being data-intensive. As a solution, ConvNeXt was introduced in 2022. It is a convolutional-based model that integrates strategies from Transformer models, resulting in an architecture that outperforms state-of-the-art Transformers while remaining data efficient. One of its key features is the separate handling of spatial and channel information, akin to the self-attention mechanism in Transformers. This property is achieved through the use of depthwise separable convolutions. We adapted the ConvNeXt architecture to our problem by utilizing a U-Net architecture with ConvNeXt as its backbone (i.e., ConvNeXt layers served as feature extractors in the encoder part). A schematic representation of the considered architecture is depicted in Figure 3.
@@ 2.4.1 Zoomed samples
Standard simulations had $h\times w$ shape, where $h$ and $w$ refer to the original number of fast-time points and number of firings, respectively. Zooming was achieved by simulating color Doppler acquisitions with a finer grid than the one used for standard simulations, and then cropping the result to an $h\times w$ shape at a random location. This was done by repeating simulations and increasing the number of fast-time points with a 1.5 ratio and the number of focused firings with a 1.5 ratio. The procedure is illustrated in Figure 4.
Figure 4: Illustration of the zooming data augmentation procedure.
@@ 2.4.2 Aliased samples
Aliasing patterns appear in Doppler velocity maps when the maximum speed exceeds the Nyquist velocity. In such cases, phase unwrapping methods can be applied to the Doppler velocity maps as a post-processing step, as described in the introduction. In contrast to these post-processing approaches, we aimed to make the Doppler estimation inherently robust to aliasing. This was achieved by providing the deep learning models with new training pairs $(S,D)$, where $S$ represents the I/Q aliased signal and $D$ denotes the corresponding alias-free Doppler ground truth. The aliased samples were obtained during simulations by randomly reducing the original PRF values of each sequence by a factor in the $[0.4,0.6]$ range. Each simulated aliased sample was considered as an augmented sample and added to the dataset.
@@ 3.1.1 In silico experiment
PRF 4500 - 7700 Hz
sector width 60°$\degree$ - 90°$\degree$
depth 10 - 17 cm
@@
During the simulations, we leveraged patient-specific parameters as pulse repetition frequency (PRF), sector angle, and depth. The simulated RF signals were demodulated into I/Q signals and beamformed into a $180\times 40$ grid using delay-and-sum. Each training pair $(S,D)$ consisted of I/Q signals $S\in\mathbb{C}^{180\times 40\times n}$ and Doppler velocity references $D\in\mathbb{R}^{180\times 40}$.
The simulated dataset is referred to as Original-Set and described in Table 2.
Table 2: Description of the three simulated datasets.
Total Non-aliased Aliased
Original-Set 2,576 2,552 24
Aliased-Set 1,142 0 1,142
Zoomed-Set 2,576 1,883 693
A sample form Original-Set is shown in Figure 5.
Figure 5: Example of a simulated echocardiography acquisition with its original and decreased PRF. The Doppler velocities were estimated using an autocorrelator in both cases.
Only 24 samples contained aliased regions, accounting for less than 0.001% of the total dataset pixels. This was circumvented with the simulation procedure described in Section 2.4.2 to increase the number of aliased samples by artificially decreasing the PRF. The set of aliased samples obtained is referred to as Aliased-Set in the sequel and is described in Table 2. An example of a sample simulated with both its original and decreased PRF is depicted in Figure 5. Finally, a third dataset was simulated to provide for augmented zoomed samples as described in Section 2.4.1. This dataset is referred to as Zoomed-Set and is described in Table 2.
From these three simulated datasets, we defined two training datasets: Train-Set-1, including Original-Set and non-aliased augmented samples of Zoomed-Set; and Train-Set-2, comprising all simulated datasets. They are described in Table 3.
Table 3: Description of the two training datasets.
Source data Total Non-aliased Aliased
Train-Set-1
Original-Set
Zoomed-Set
4,459 4,435 24
Train-Set-2
Original-Set
Aliased-Set
Zoomed-Set
6,294 4,435 1,859
Finally, samples with reduced packet size were created by selecting two consecutive frames in the slow-time.
@@ 3.1.2 In vitro experiment
We designed a rotating disk experiment to measure Doppler velocities from a color Doppler acquisition. We constructed a 6 cm diameter agar disk and connected it to a rotating motor. A Verasonics P4-2v cardiac probe was positioned at 8 cm from the center of the disk. To minimize clutter and artifacts, we submerged the entire setup in a water tank.
The acquisition parameters were set as described in Table 1, with the sector width, depth and PRF configured to $50^{\circ}$, 12 cm and 6,000 Hz, respectively. The packet size was $n=32$. We ran a total of 12 experiments, varying the maximum outer speed between 0.1 and 1.13 m/s, mirroring intracardiac blood velocities. The Nyquist velocity was $v_{N}=0.85$ m/s for all experiments, resulting in four experiments displaying aliased regions.
@@ 3.1.3 In vivo experiment
Cardiac echo-Doppler focused acquisitions were conducted on six healthy volunteers. The study received approval from the ethics and research committee of Polytechnique Montréal (CER-2122-54-D). In accordance with the simulation and phantom experiments, the P4-2v Verasonics probe was used with identical parameters as those used for the phantom. Data were acquired in the apical three-chamber view, including both B-mode and Doppler information. The packet size was set to $n=16$. The ultrasound signals were post-processed and beamformed on a $180\times 40$ grid. We obtained 8 to 16 color Doppler image samples from each volunteer.
To be consistent with the clutter-free setting of both in silico and in vitro experiments, we applied an SVD-based clutter filter. We discarded the first four largest singular values to isolate the blood I/Q signals. We segmented the left ventricular endocardium to mask off I/Q signals not pertaining to intraventricular blood. Samples with low Doppler power in over 70% of the ventricular region were discarded, resulting in 50 selected samples. Only one of the selected samples contained aliased regions. To assess the method’s robustness, we deliberately introduced aliasing by considering one out of two (or three) frames in the slow-time, thereby artificially decreasing the PRF by a factor of two (or three). We applied this procedure to the 50 selected samples, providing 20 new samples with several aliased regions.
Samples with reduced packet size were generated by selecting two consecutive frames in the slow-time. In the in vivo experiments, ground truth Doppler velocity maps were not available. Hence, Doppler velocity maps obtained with the autocorrelator using a packet size of $n=8$ were used as references.
@@ 3.2.1 Training settings
Because the ultrasound data were simulated with patient-specific PRFs, the Nyquist velocity varied from sample to sample. Therefore, we opted to use phase shifts as the ground truth instead of Doppler velocities to ensure standardized inputs across samples. Each sample was normalized with respect to its maximum complex modulus across the slow-time. For greater robustness, we trained the network using 9-fold cross-validation, where samples from a given patient were assigned to the same fold. Once the test fold was chosen, the split of samples in training/validation was set to 9/1 while ensuring that samples from a given patient remained in the same split. The network weights were initialized with the Xavier distribution. The loss function was the masked mean squared error, where the mask corresponded to the left ventricular cavity. The optimizer was AdamW, and the batch size was set to 16. The initial value of the learning rate was set to $10^{-3}$ and decreased following a plateau scheduler with a patience of 10 epochs. Trainings were performed using the PyTorch library on an NVIDIA V100 GPU with 16 GB of memory.
@@ 4.1 In silico results
The performance of our methods in Doppler velocity estimation, when trained on Train-Set-1 and using a packet size $n=2$, is presented in Table 4. For all methods, the metrics were computed exclusively within the left ventricle.
Table 4: Doppler velocity estimation results of the baseline method and the three proposed deep learning methods, trained on Train-Set-1 and evaluated on Original-Set, Aliased-Set and aliased pixels of Aliased-Set. The results correspond to the mean ± standard deviation of the RMSE [cm/s].
Method Original-Set Aliased-Set Aliased pixels
Autocorrelator $3.7\pm 2.6$ $16\pm 9.8$ $57\pm 23$
Real U-Net $1.8\pm 0.9$ $5.4\pm 5.3$ $14\pm 14$
Complex U-Net $1.8\pm 0.8$ $\boldsymbol{5.3\pm 5.1}$ $\boldsymbol{14\pm 13}$
ConvNeXt $\boldsymbol{1.7\pm 1.1}$ $6.1\pm 6.3$ $16\pm 15$
@@
We examined the samples with higher errors for both the autocorrelator and the deep learning models. For the former, large estimation errors mainly occurred in two situations: noisy signals and in the presence of aliasing. For the deep learning models, large errors were observed in two specific settings: in a few pixels near the valves and the presence of aliasing across a large region of the ventricle. To mitigate the effects of aliasing, we trained our models on Train-Set-2. The results of the estimation are presented in Table 5.
Table 5: Doppler velocity estimation results of the baseline method and the three proposed deep learning methods trained on Train-Set-2 and evaluated on Original-Set, Aliased-Set and aliased pixels of Aliased-Set. The results represent the mean ± standard deviation of the RMSE [cm/s].
Method Original-Set Aliased-Set Aliased pixels
Autocorrelator $3.7\pm 2.6$ $16\pm 9.8$ $57\pm 23$
Real U-Net $2.0\pm 0.8$ $2.4\pm 1.1$ $3.7\pm 3.1$
Complex U-Net $2.0\pm 0.8$ $2.5\pm 1.3$ $3.7\pm 3.0$
ConvNeXt $\boldsymbol{1.8\pm 0.7}$ $\boldsymbol{2.1\pm 1.0}$ $\boldsymbol{3.3\pm 2.4}$
We noted that when testing on Original-Set, we maintained a consistent performance compared to the models trained on Train-Set-1. Moreover, when testing on Aliased-Set, we observed a substantial reduction in error compared to the autocorrelator method, indicating that the models are capable to alleviate aliasing to some extent.
@@
Figure 6: Doppler velocity estimations for four in silico samples. B-mode and color Doppler images were simulated from the data of four different patients. Samples A and B belonged to Original-Set and had velocities below the Nyquist limit. Samples C and D had velocities belonged to Aliased-Set above the Nyquist limit. Aliased regions are indicated with white arrows.
We report predictions obtained with ConvNeXt (the best-performing model) and Complex U-Net. The Real U-Net provided results similar to the Complex U-Net. As explained above, our models trained with both Train-Set-1 and Train-Set-2 achieved accurate estimates in the absence of aliasing (Sample A and Sample B). In such cases, we observed that the autocorrelator was more susceptible to noise. When encountering small aliased regions (Sample C), the models trained with Train-Set-1 either effectively removed aliasing (see the Complex U-Net estimation), or tended to smooth the region, leading to velocity underestimation (see the right part of the ConvNeXt estimation). When trained with Train-Set-2, the estimation was closer to the ground truth (see the ConvNeXt estimation). In the case of larger aliased regions as in Sample D, the models trained on Train-Set-1 systematically smoothed the aliased region and hence underestimated velocities, whereas when trained with Train-Set-2 they yielded results closely aligned with the ground truth.